OpenStudy (anonymous):

What expression equals to 2√8m^2n^3

OpenStudy (driftracer305):

\[2\sqrt{8m^2} n^2\]

OpenStudy (driftracer305):

is that it or is n^3 inside the radical sign?

OpenStudy (anonymous):

yes

OpenStudy (driftracer305):

so it should be 4m sqrt 2 n^3

OpenStudy (driftracer305):

\[4m \sqrt{2} n^3\]

OpenStudy (driftracer305):

so wait i didnt quite get u @triciav .... is um n^3 under the sqrt sign??

OpenStudy (anonymous):

yeah it is

OpenStudy (driftracer305):

ohh ok

OpenStudy (driftracer305):

so it should be \[4m n^(3/2) \sqrt{2}\]

OpenStudy (anonymous):

Really? this isn't on the answer choices

OpenStudy (driftracer305):

can i see the choices?

OpenStudy (anonymous):

\[4m ^{2}n \sqrt{2n} \]

OpenStudy (anonymous):

\[4m ^{3}n ^{2}\sqrt{2n}\]

OpenStudy (anonymous):

\[4m ^{2}n \sqrt{2n}\]

OpenStudy (anonymous):

\[4mn \sqrt{2n}\]

OpenStudy (anonymous):

those are the 4 answer choices

OpenStudy (driftracer305):

yup..... its the last one..... u culd do n^3/2 or do n sqrt 2n..... both work

OpenStudy (driftracer305):

so D.......... :D

OpenStudy (anonymous):

thank you!!! :D

OpenStudy (driftracer305):

sure..... no problem....